# Other Vehicle Trajectories Are Also Needed: A Driving World Model Unifies Ego-Other Vehicle Trajectories in Video Latent Space

### v3

@@ 1 Introduction
End-to-end autonomous driving [20, 33, 26] has gained increasing attention recently, since the approach integrates all modules (e.g. perception, decision, and planning modules) into a model optimized jointly to directly output planning results based on input multi-sensor data. Despite the promising performance of the end-to-end autonomous driving model, effectively handling out-of-distribution scenarios continues to be a significant challenge, particularly since such situations are often hazardous and costly to simulate. The world model [38, 39, 9, 15] can predict future driving scenes based on historical observations and future driving actions, which is a potential solution to evaluate the autonomous driving model and avoid catastrophic errors.
@@
Some world models [9, 42] explore predicting future driving scenes with the trajectory, since it can reflect the driving action of the vehicle in the scene more precisely. Despite advanced end-to-end autonomous driving systems (e.g. VAD [21]) can predict other vehicles’ motions and plan ego vehicle’s trajectory, these world models primarily focus on the ego vehicle trajectory. They view the trajectory as a series of points in the bird’s eye view (BEV) space [19, 25], which are directly encoded as the condition for video generation. However, there are three main drawbacks in the approaches mentioned above. Firstly, they just consider the ego vehicle trajectory and leave other vehicles uncontrollable in the generated video. As a result, the model cannot realistically simulate the interaction between the ego vehicle and the driving scenario, and generate diverse novel scenes by changing other vehicle trajectories as well. Secondly, the distribution of the encoding based on the trajectory points in the BEV space is quite different from that of the video latents in visual modality without aligning their feature space. In addition, it is impractical to correspond multiple trajectories in the BEV space to numerous vehicles in visual modality, since the BEV space is mismatched with the video space.
@@
To tackle above issue, a driving world model unifying ego-other vehicle trajectories in videos named EOT-WM is proposed in this paper. As is shown in Fig. 1, Vista [9] only uses the ego vehicle trajectory, and the stationary other vehicle in the groundtruth video also moves forward in the generated video. The proposed EOT-WM can generate more realistic videos with controllable ego and other vehicle actions. In addition, the model can also generate video matched based on novel self-produced trajectories. Specifically, instead of representing the trajectory via points in the BEV space, we project these points into the image coordinate and plot the ego vehicle trajectory and the other vehicle trajectories in separate blank videos to generate trajectory videos for learning in unified visual modality. Then, we adopt Spatial-Temporal Variational Auto Encoder (STVAE) to encode the scene video and trajectory video to achieve scene video latents and trajectory latents with sharing feature space. Moreover, the scene and trajectory latents achieved in this manner are aligned temporally and spatially to realize effective control. Finally, we design Trajectory-injected Diffusion Transformer (TiDiT) to integrate motion guidances provided by trajectory latents into video latents for denoising the noisy video latents more precisely. As a result, the entire model can predict future frames based on the given initial frames with the control of text and trajectory. To evaluate the controllability of trajectories, we propose a metric based on control latent similarity, which compares the predicted trajectory latents with the groundtruth trajectory latents.
@@
  • We conduct comprehensive experiments on the nuScenes dataset to evaluate the proposed EOT-WM, which outperforms the state-of-the-art method by 30% in FID and 55% in FVD.
2 Related Work
2.1 Video Generation
@@
Video generation is capable of understanding the world and generating realistic video samples. Various kinds of models have been studied in the past, including VAE-based [35, 8], flow-based [7, 24], GAN-based [4, 44] and auto-regressive models [10, 41]. Recently, diffusion models have achieved breakthroughs in image generation [30, 31], and diffusion models are also introduced into video generation [3, 11, 18, 43]. Most of these models [3, 11, 18, 43] use the text as the condition to control video generation. To explore more condition forms for controlling video generation, image-to-video generative models [2, 6] are studied since the image can provide more specific priors. However, above models [3, 11, 18, 43, 2, 6] are incapable of predicting the future state, which are crucial in autonomous driving. In addition, camera motion [40] and object motion [40, 45] are employed to realize more flexible video generation. Unfortunately, these models are not specifically designed to tackle complicated autonomous driving videos and cannot model vehicle actions well. In this paper, we propose EOT-WM that jointly uses texts and trajectories to predict future video frames based on given initial states.
Figure 2: Illustration of the proposed EOT-WM, which consists of three key parts: (1) Video-based Trajectory Representation, (2) Aligned Motion Guidance Generation, (3) Trajectory-injected Diffusion Transformer.
2.2 World Model
@@
The world model can predict the future states based on historical observations and future actions, which has been extensively applied in simulated games [12, 13, 14] and indoor embodiment [22, 36, 29]. Recently, the world model has attracted great attention in autonomous driving [38, 28, 39, 42, 9, 15, 1]. WoVoGen [28] and Drive-WM [39] are devoted to generating 6-view driving videos based on the given conditions. Drivedreamer [38], GenAD [42], Vista [9], and GEM [15] develop world models to predict future driving scenes via the front-view video. However, most world models mentioned above concentrate on the actions of the ego vehicle and ignore those of the other vehicles, making the interaction with the environment insufficient in these world models. Although GEM [15] uses future object features and human poses to generate future videos, it is more impractical for autonomous driving systems to obtain them compared with future trajectories. Particularly, Cosmos [1] is a commercial model comprising over 7 billion parameters, which imposes substantial computational overhead during both training and inference. In this paper, the proposed EOT-WM uses the ego-other vehicle trajectories to represent the motions of vehicles in the scene, and generate more realistic videos with all vehicles controllable.
3 EOT-WM Framework
@@
The overall architecture of the proposed EOT-WM framework is illustrated in Fig. 2. We build EOT-WM based on CogvideoX [43], and modify it as the world model architecture with injected trajectory condition. Specifically, the proposed EOT-WM framework consists of Video-based Trajectory Representation (VTR), Aligned Motion Guidance Generation (AMGG), and Trajectory-injected Diffusion Transformer (TiDiT), which are described in detail in the following sections.
3.1 Video-based Trajectory Representation
@@
Figure 3: Illustration of the original video, other vehicle trajectory (O. V. Traj.) and ego vehicle trajectory (E. V. Traj.) used for the proposed EOT-WM. To be brief, we only visualize the 1st, 7th, 13th, 19th, 25th frames in the figure.
@@
The trajectory in the end-to-end autonomous driving system is usually represented as a series of points in the BEV space for planning, which can be formulated as $\text{Traj}^{BEV}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{T},y_{T})\}$ and $T$ is the number of points. However, the world model in this paper aims to generate a video that simulates the driving scene based on vehicle trajectories. All elements in the generated video such as vehicles and roads are represented by pixels, leading to the natural mismatch between the points in the BEV space and the pixels in the video. In addition, it is impractical to correspond multiple trajectories in the BEV space to numerous vehicles in the video. Therefore, we propose to represent the trajectory in the form of video to act as the condition for the world model.
@@
where $R$ and $B$ are rotation matrix and translation matrix of the camera extrinsic parameters for the first frame, and $K$ is the the camera intrinsic parameters. Then, as is shown in Fig. 3, we plot the ego vehicle trajectory and the other vehicle trajectories in separate blank videos to generate trajectory videos. For frame $t$ in the trajectory video, only current and future points $\text{Traj}^{plot}=\{(x_{t}^{I},y_{t}^{I}),...,(x_{T}^{I},y_{T}^{I})\}$ are plotted, indicating the future motion of the vehicle. As a result, the trajectory video contains the motion information of the vehicle and corresponds each trajectory to the vehicle in the video.
3.2 Aligned Motion Guidance Generation
After obtaining trajectory videos, the trajectory can be learned in unified visual modality with driving scene videos. The world model needs to encode trajectory videos as conditions to guide video generation. Here, we adopt the Spatial-Temporal Variational Auto Encoder (STVAE) in CogvideoX [43] to encode them. On the one hand, STVAE is capable of extracting high-quality spatial-temporal features of trajectory videos, which are aligned with driving video latents temporally and spatially. On the other, the condition features can share the feature space with the driving video latents since they both use the same STVAE to generate latents, which is easier to be learned jointly. The vehicle trajectory latents $z_{traj}$ and original video latents $z_{vid}$ for training can be formulated as follows:
@@ 3.2 Aligned Motion Guidance Generation
where $V_{traj}$, $V_{ori}$ are the ego vehicle trajectory video and origin driving video, respectively. $\Psi$ denotes the STVAE in the model. $z_{traj}$ is used to provide motion guidance for video generation.
3.3 Trajectory-injected Diffusion Transformer
Typical Diffusion Transformer in CogvideoX [43] aims to predict the noise $n^{a}$ added to $z_{vid}$ based on the noisy video latent $z_{vid}^{noise}$, which is formulated as
@@ 3.3 Trajectory-injected Diffusion Transformer
In the inference stage, Denoising Diffusion Probabilistic Model (DDPM) [17] can denoise from random sampling noise to video latents for the VAE decoder to generate video. However, the video cannot be generated with specific initial frames in such a manner, which is important for the world model. In addition, the original CogvideoX [43] can only use the text as the condition to control video generation.
@@
To further inject trajectory conditions, frame latents $z$ and motion guidance $z_{traj}$ sharing the feature space are concatenated in channel dimension. As a result, the video and condition latents are fused and aligned both spatially and temporally. Then, 3D convolution layers are adopted for learning to obtain the final visual latents $z_{vid}^{f}$, which can be calculated as
@@
The original Diffusion Transformer in CogvideoX [43] uses the text as the condition to control video generation. To fully reserve the ability of CogvideoX, we utilize the scene caption provided in OmniDrive [37] to train the proposed TiDiT. As a result, visual latents $z_{vid}^{f}$ and text embedding $z_{text}$ extracted from the scene caption are concatenated and fed into several layers of Expert Diffusion Transformer in CogvideoX to predict the noise added to $z_{vid}$ as follows:
@@
where $\Omega$ represents the MLP. We design a control latent similarity loss $l_{traj}$ to train the MLP, and the loss is formulated as
@@
4 Experiments
4.1 Dataset and Evaluation Metric
Extensive experiments are conducted on the nuScenes dataset [5] to evaluate the proposed EOT-WM. Following the setup of Vista [9] in the nuScenes dataset, the training and evaluation set contain 25109 and 5369 videos with 25 frames. The ego vehicle trajectory corresponding to the video can be directly obtained in the annotations. The other vehicle trajectory is only annotated in key frames, and we use interpolation based on the annotated trajectories to complete the full trajectory corresponding to the 25-frame video. We use FID [16] and FVD [34] to evaluate the quality of generated videos. Moreover, we provide qualitative demonstrations that underscore the advancements of the proposed EOT-WM.
@@ 4.1 Dataset and Evaluation Metric
4.2 Implementation Details
We build the proposed EOT-WM based on CogvideoX-2B [43]. We use the first video latent as the context for predicting the future frames. The frame rate of the generated video is 10Hz. We use the AdamW [27] optimizer with a learning rate of $2\times 10^{-5}$ to train the model. By default, the proposed model is trained at 768×1280 resolution for 60 epochs with 64 NVIDIA A800 GPUs with a total batch size of 128.
4.3 Comparison with State-of-the-art Methods
Quantitative Result. Comparison results of FID and FVD for video generation achieved by the proposed EOT-WM and competing methods on the validation set of nuScenes dataset are given in Tab. 1. As can be observed, the state-of-the-art driving world model Vista [9] can generate videos with higher quality and resolution compared with previous works [32, 38, 28, 39, 42]. However, Vista [9] cannot control other vehicles in the scene, and using trajectories in BEV space is not sufficiently effective. Therefore, the proposed EOT-WM controlling video generation with ego and other vehicle trajectories in video space achieves 4.8 FID and 40.0 FVD, surpassing Vista by 30% in FID and 55% in FVD. Moreover, the proposed EOT-WM can generate videos in 768×1280 resolution, superior to previous world models as well.
Figure 5: Instance of difference text types used for the proposed EOT-WM, where the scene description and action description are provided in OmniDrive [37].
@@ 4.3 Comparison with State-of-the-art Methods
Methods Resolution FID $\downarrow$ FVD $\downarrow$ DriveGAN [32] 256×256 73.4 502.3 DriveDreamer [38] 128×192 52.6 452.0 WoVoGen [28] 256×448 27.6 417.7 Drive-WM [39] 192×384 15.8 122.7 GenAD [42] 256×448 15.4 184.0 Vista [9] 576×1024 6.9 89.4 EOT-WM 768×1280 4.8 40.0
@@
Qualitative Result. To further demonstrate the superiority in action controllability of the proposed EOT-WM, we provide several representative cases achieved by the proposed EOT-WM compared with Vista [9] in Fig. 4. As is shown in Fig. 4 (a), the blank car in the red rectangle stands by the side of the road. Vista and EOT-WM w/o O.V. Traj. do not consider the other vehicle trajectories, and the blank car moves forward in the generated videos. In contrast, the blank car keeps still in the video generated by the proposed EOT-WM. In Fig. 4 (b), the truck is backing up on the road. However, Vista and EOT-WM w/o O.V. Traj. both generate the truck moving forward in the generated video. As for the proposed EOT-WM, the action of the truck is predicted correctly in the generated video. These cases demonstrate that the proposed EOT-WM can generate more realistic videos with ego and other vehicle trajectories.
@@
Table 2: Ablation for key components on the validation set of nuScenes dataset.
4.4 Ablation Study
@@
4.5 Effect of difference text types
The original CogvideoX [43] is trained to use the text to control video generation, and other advanced video generation models [11, 23] are also text-to-video models. When building the world model based on such text-to-video models, it is unadvisable to remove the text element since the capability of the original text-to-video model will be damaged. As can be observed in Tab. 3, while the proposed EOT-WM does not use the text, FID and FVD achieved decline obviously. Then, we adopt three types of texts to control the video generation model. The performances obtained by these types of texts are similar, demonstrating that the video generation model can be trained to adapt different types of texts. Among them, EOT-WM using the action description achieves the best performance since the caption contains action information. However, considering the capability of generating novel scenes, we choose the scene description for EOT-WM as the action description will conflict with the novel trajectories.
4.6 Generating novel scenes
@@
4.7 Evaluating Trajectory Controllability
@@
5 Conclusion
In this work, a novel driving world model named EOT-WM is proposed, which can use ego and other vehicle trajectories to generate realistic videos. Specifically, we propose to represent ego and other vehicle trajectories in the video space instead of the BEV space for learning in the unified visual modality. Then, AMGG is proposed to generate trajectory latents aligned with driving video latents spatially and temporally, which also share feature space. Finally, TiDiT is designed to denoise the noisy video latents more precisely via jointly using the motion guidances based on the ego-other vehicle trajectories. Experiments conducted on the nuScenes dataset demonstrate the superiority of the proposed EOT-WM.
In the future work, we will explore the manner to control the other vehicles more precisely and the quantitative metric to evaluate the consistency between the given trajectory and the generated video.

### v4

@@ Introduction
End-to-end autonomous driving (Hu et al. 2023; Tian et al. 2024; Li et al. 2025) has gained increasing attention recently, since the approach integrates all modules (e.g. perception, decision, and planning modules) into a model optimized jointly to directly output planning results based on input multi-sensor data. Despite the promising performance of the end-to-end autonomous driving model, effectively handling out-of-distribution scenarios continues to be a significant challenge, particularly since such situations are often hazardous and costly to simulate. The world model (Wang et al. 2023b, 2024b; Gao et al. 2024; Hassan et al. 2025) can predict future driving scenes based on historical observations and future driving actions, which is a potential solution to evaluate the autonomous driving model and avoid catastrophic errors.
@@
Some world models (Gao et al. 2024; Yang et al. 2024a) explore predicting future driving scenes with the trajectory, since it can reflect the driving action of the vehicle in the scene more precisely. Despite advanced end-to-end autonomous driving systems (e.g. VAD (Jiang et al. 2023)) can predict other vehicles’ motions and plan ego vehicle’s trajectory, these world models primarily focus on the ego vehicle trajectory. They view the trajectory as a series of points in the bird’s eye view (BEV) space (Hu et al. 2021; Li et al. 2024), which are directly encoded as the condition for video generation. However, there are three main drawbacks in the approaches mentioned above. Firstly, they just consider the ego vehicle trajectory and leave other vehicles uncontrollable in the generated video. As a result, the model cannot realistically simulate the interaction between the ego vehicle and the driving scenario, and generate diverse novel scenes by changing other vehicle trajectories as well. Secondly, the distribution of the encoding based on the trajectory points in the BEV space is quite different from that of the video latents in visual modality without aligning their feature space. In addition, it is impractical to correspond multiple trajectories in the BEV space to numerous vehicles in visual modality, since the BEV space is mismatched with the video space.
@@
To tackle above issue, a driving world model unifying ego-other vehicle trajectories in videos named EOT-WM is proposed in this paper. As is shown in Fig. 1, Vista (Gao et al. 2024) only uses the ego vehicle trajectory, and the stationary other vehicle in the groundtruth video also moves forward in the generated video. The proposed EOT-WM can generate more realistic videos with controllable ego and other vehicle actions. In addition, the model can also generate video matched based on novel self-produced trajectories. Specifically, instead of representing the trajectory via points in the BEV space, we project these points into the image coordinate and plot the ego vehicle trajectory and the other vehicle trajectories in separate blank videos to generate trajectory videos for learning in unified visual modality. Then, we adopt Spatial-Temporal Variational Auto Encoder (STVAE) to encode the scene video and trajectory video to achieve scene video latents and trajectory latents with sharing feature space. Moreover, the scene and trajectory latents achieved in this manner are aligned temporally and spatially to realize effective control. Finally, we design Trajectory-injected Diffusion Transformer (TiDiT) to integrate motion guidances provided by trajectory latents into video latents for denoising the noisy video latents more precisely. As a result, the entire model can predict future frames based on the given initial frames with the control of text and trajectory. To evaluate the controllability of trajectories, we propose a metric based on control latent similarity, which compares the predicted trajectory latents with the groundtruth trajectory latents.
@@
Related Work
Video Generation
@@
Video generation is capable of understanding the world and generating realistic video samples. Various kinds of models have been studied in the past, including VAE-based (Villegas et al. 2019; Franceschi et al. 2020), flow-based (Dorkenwald et al. 2021; Kumar et al. 2020), GAN-based (Brooks et al. 2022; Yu et al. 2022) and auto-regressive models (Ge et al. 2022; Weissenborn, Täckström, and Uszkoreit 2020). Recently, diffusion models have achieved breakthroughs in image generation (Nichol et al. 2022; Rombach et al. 2022), and diffusion models are also introduced into video generation (Blattmann et al. 2023b; Guo et al. 2024; Ho et al. 2022; Yang et al. 2024b). Most of these models (Blattmann et al. 2023b; Guo et al. 2024; Ho et al. 2022; Yang et al. 2024b) use the text as the condition to control video generation. To explore more condition forms for controlling video generation, image-to-video generative models (Blattmann et al. 2023a; Chen et al. 2023) are studied since the image can provide more specific priors. However, above models (Blattmann et al. 2023b; Guo et al. 2024; Ho et al. 2022; Yang et al. 2024b; Blattmann et al. 2023a; Chen et al. 2023) are incapable of predicting the future state, which are crucial in autonomous driving. In addition, camera motion (Wang et al. 2024c) and object motion (Wang et al. 2024c; Zhang et al. 2024) are employed to realize more flexible video generation. Unfortunately, these models are not specifically designed to tackle complicated autonomous driving videos and cannot model vehicle actions well. In this paper, we propose EOT-WM that jointly uses texts and trajectories to predict future video frames based on given initial states.
Figure 2: Illustration of the proposed EOT-WM.
World Model
@@
The world model can predict the future states based on historical observations and future actions, which has been extensively applied in simulated games (Ha and Schmidhuber 2018; Hafner et al. 2020, 2021) and indoor embodiment (Koh et al. 2021; Wang et al. 2023a; Mendonca, Bahl, and Pathak 2023). Recently, the world model has attracted great attention in autonomous driving (Wang et al. 2023b; Lu et al. 2024; Wang et al. 2024b; Yang et al. 2024a; Gao et al. 2024; Hassan et al. 2025; Agarwal et al. 2025). WoVoGen (Lu et al. 2024) and Drive-WM (Wang et al. 2024b) are devoted to generating 6-view driving videos based on the given conditions. Drivedreamer (Wang et al. 2023b), GenAD (Yang et al. 2024a), Vista (Gao et al. 2024), and GEM (Hassan et al. 2025) develop world models to predict future driving scenes via the front-view video. However, most world models mentioned above concentrate on the actions of the ego vehicle and ignore those of the other vehicles, making the interaction with the environment insufficient in these world models. Although GEM (Hassan et al. 2025) uses future object features and human poses to generate future videos, it is more impractical for autonomous driving systems to obtain them compared with future trajectories. Particularly, Cosmos (Agarwal et al. 2025) is a commercial model comprising over 7 billion parameters, which imposes substantial computational overhead during both training and inference. In this paper, the proposed EOT-WM uses the ego-other vehicle trajectories to represent the motions of vehicles in the scene, and generate more realistic videos with all vehicles controllable.
EOT-WM Framework
@@
The overall architecture of the proposed EOT-WM framework is illustrated in Fig. 2. We build EOT-WM based on CogvideoX (Yang et al. 2024b), and modify it as the world model architecture with injected trajectory condition. Specifically, the proposed EOT-WM framework consists of Video-based Trajectory Representation (VTR), Aligned Motion Guidance Generation (AMGG), and Trajectory-injected Diffusion Transformer (TiDiT), which are described in detail in the following sections.
@@
Figure 3: Illustration of the original video, other vehicle trajectory (O. V. Traj.) and ego vehicle trajectory (E. V. Traj.) used for the proposed EOT-WM. To be brief, we only visualize the 1st, 7th, 13th, 19th, 25th frames.
Video-based Trajectory Representation
@@
The trajectory in the end-to-end autonomous driving system is usually represented as a series of points in the BEV space for planning, which can be formulated as $\text{Traj}^{BEV}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{T},y_{T})\}$ and $T$ is the number of points. However, the world model in this paper aims to generate a video that simulates the driving scene based on vehicle trajectories. All elements in the generated video such as vehicles and roads are represented by pixels, leading to the natural mismatch between the points in the BEV space and the pixels in the video. In addition, it is impractical to correspond multiple trajectories in the BEV space to numerous vehicles in the video. Therefore, we propose to represent the trajectory in the form of video to act as the condition.
@@
where $R$ and $B$ are rotation matrix and translation matrix of the camera extrinsic parameters for the first frame, and $K$ is the camera intrinsic parameters. Then, as is shown in Fig. 3, we plot the ego vehicle trajectory and other vehicle trajectories in separate blank videos to generate trajectory videos. For frame $t$ in the trajectory video, only current and future points $\text{Traj}^{plot}=\{(x_{t}^{I},y_{t}^{I}),...,(x_{T}^{I},y_{T}^{I})\}$ are plotted, indicating the future motion of the vehicle. As a result, the trajectory video contains the motion information of vehicles and corresponds each trajectory to the vehicle in the video.
Aligned Motion Guidance Generation
After obtaining trajectory videos, the trajectory can be learned in unified visual modality with driving scene videos. The world model needs to encode trajectory videos as conditions to guide video generation. Here, we adopt the Spatial-Temporal Variational Auto Encoder (STVAE) in CogvideoX (Yang et al. 2024b) to encode them. On the one hand, STVAE is capable of extracting high-quality spatial-temporal features of trajectory videos, which are aligned with driving video latents temporally and spatially. On the other, the condition features can share the feature space with the driving video latents since they both use the same STVAE to generate latents, which is easier to be learned jointly. The vehicle trajectory latents $z_{traj}$ and original video latents $z_{vid}$ for training can be formulated as follows:
@@
where $V_{traj}$, $V_{ori}$ are the ego vehicle trajectory video and original driving video, respectively. $\Psi$ denotes the STVAE in the model. $z_{traj}$ is used to provide motion guidance for video generation.
Trajectory-injected Diffusion Transformer
Typical Diffusion Transformer in CogvideoX (Yang et al. 2024b) aims to predict the noise $n^{a}$ added to $z_{vid}$ based on the noisy video latent $z_{vid}^{noise}$, which is formulated as
@@
In the inference stage, Denoising Diffusion Probabilistic Model (DDPM) (Ho, Jain, and Abbeel 2020) can denoise from random sampling noise to video latents for the VAE decoder to generate video. However, the video cannot be generated with specific initial frames in such a manner, which is important for the world model. In addition, the original CogvideoX (Yang et al. 2024b) can only use the text as the condition to control video generation.
@@
To further inject trajectory conditions, frame latents $z$ and motion guidance $z_{traj}$ sharing the feature space are concatenated in channel dimension. As a result, the video and condition latents are fused and aligned spatially and temporally. Then, 3D convolution layers are adopted for learning to obtain the final visual latents $z_{vid}^{f}$, which can be calculated as
@@
The original Diffusion Transformer in CogvideoX (Yang et al. 2024b) uses the text as the condition to control video generation. To fully reserve the ability of CogvideoX, we utilize the scene caption provided in OmniDrive (Wang et al. 2024a) to train the proposed TiDiT. As a result, visual latents $z_{vid}^{f}$ and text embedding $z_{text}$ extracted from the scene caption are concatenated and fed into several layers of Expert Diffusion Transformer in CogvideoX to predict the noise added to $z_{vid}$ as follows:
@@
where $\Omega$ represents the MLP. We design a control latent similarity loss $l_{traj}$ to train the MLP, which is formulated as
@@
Experiments
Dataset and Evaluation Metric
Extensive experiments are conducted on the nuScenes dataset (Caesar et al. 2020) to evaluate the proposed EOT-WM. Following the setup of Vista (Gao et al. 2024) in the nuScenes dataset, the training and evaluation set contain 25109 and 5369 videos with 25 frames. The ego vehicle trajectory corresponding to the video can be directly obtained in the annotations. The other vehicle trajectory is only annotated in key frames, and we use interpolation based on the annotated trajectories to complete the full trajectory corresponding to the 25-frame video. We use FID (Heusel et al. 2017) and FVD (Unterthiner et al. 2018) to evaluate the quality of generated videos. Moreover, we provide qualitative demonstrations that underscore the advancements of the proposed EOT-WM.
Implementation Details
We build the proposed EOT-WM based on CogvideoX-2B (Yang et al. 2024b). We use the first video latent as the context for predicting the future frames. The frame rate of the generated video is 10Hz. We use the AdamW (Loshchilov 2017) optimizer with a learning rate of $2\times 10^{-5}$ to train the model. By default, the proposed model is trained at 768×1280 resolution for 60 epochs with 64 NVIDIA A800 GPUs with a total batch size of 128.
Comparison with State-of-the-art Methods
Quantitative Result. Comparison results of FID and FVD for video generation achieved by the proposed EOT-WM and competing methods on the validation set of nuScenes dataset are given in Tab. 1. As can be observed, the state-of-the-art driving world model Vista (Gao et al. 2024) can generate videos with higher quality and resolution compared with previous works (Santana and Hotz 2016; Wang et al. 2023b; Lu et al. 2024; Wang et al. 2024b; Yang et al. 2024a). However, Vista (Gao et al. 2024) cannot control other vehicles in the scene, and using trajectories in BEV space is not sufficiently effective. Therefore, the proposed EOT-WM controlling video generation with ego and other vehicle trajectories in video space achieves 4.8 FID and 40.0 FVD, surpassing Vista by 30% in FID and 55% in FVD. Moreover, the proposed EOT-WM can generate videos in 768×1280 resolution, superior to previous world models as well.
Figure 5: Instance of difference text types used for the proposed EOT-WM, where the scene description and action description are provided in OmniDrive (Wang et al. 2024a).
@@
Methods Resolution FID $\downarrow$ FVD $\downarrow$ DriveGAN (Santana and Hotz 2016) 256×256 73.4 502.3 DriveDreamer (Wang et al. 2023b) 128×192 52.6 452.0 WoVoGen (Lu et al. 2024) 256×448 27.6 417.7 Drive-WM (Wang et al. 2024b) 192×384 15.8 122.7 GenAD (Yang et al. 2024a) 256×448 15.4 184.0 Vista (Gao et al. 2024) 576×1024 6.9 89.4 EOT-WM 768×1280 4.8 40.0
@@
Table 2: Ablation for key components on the validation set.
Qualitative Result. To further demonstrate the superiority in action controllability of the proposed EOT-WM, we provide several representative cases achieved by the proposed EOT-WM compared with Vista (Gao et al. 2024) in Fig. 4. As is shown in Fig. 4 (a), the blank car in the red rectangle stands by the side of the road. Vista and EOT-WM w/o O.V. Traj. do not consider the other vehicle trajectories, and the blank car moves forward in the generated videos. In contrast, the blank car keeps still in the video generated by the proposed EOT-WM. In Fig. 4 (b), the truck is backing up on the road. However, Vista and EOT-WM w/o O.V. Traj. both generate the truck moving forward in the generated video. As for the proposed EOT-WM, the action of the truck is predicted correctly in the generated video. These cases demonstrate that the proposed EOT-WM can generate more realistic videos with ego and other vehicle trajectories.
Ablation Study
@@
Effect of difference text types
The original CogvideoX (Yang et al. 2024b) is trained to use the text to control video generation, and other advanced video generation models (Guo et al. 2024; Kong et al. 2024) are also text-to-video models. When building the world model based on such text-to-video models, it is unadvisable to remove the text element since the capability of the original text-to-video model will be damaged. As can be observed in Tab. 3, while the proposed EOT-WM does not use the text, FID and FVD achieved decline obviously. Then, we adopt three types of texts to control the video generation model. The performances obtained by these types of texts are similar, demonstrating that the video generation model can be trained to adapt different types of texts. Among them, EOT-WM using the action description achieves the best performance since the caption contains action information. However, considering the capability of generating novel scenes, we choose the scene description for EOT-WM as the action description will conflict with the novel trajectories.
Generating novel scenes
@@
Evaluating Trajectory Controllability
@@
Conclusion
In this work, a novel driving world model named EOT-WM is proposed, which can use ego and other vehicle trajectories to generate realistic videos. Specifically, we propose to represent ego and other vehicle trajectories in the video space instead of the BEV space for learning in the unified visual modality. Then, AMGG is proposed to generate trajectory latents aligned with driving video latents spatially and temporally, which also share feature space. Finally, TiDiT is designed to denoise the noisy video latents more precisely via jointly using the motion guidances based on the ego-other vehicle trajectories. Experiments conducted on the nuScenes dataset demonstrate the superiority of the proposed EOT-WM. In the future work, we will explore the manner to control the other vehicles more precisely and the quantitative metric to evaluate the consistency between the given trajectory and the generated video.